OpenStudy (anonymous):

Solve for g. -6g + 3g + 12 > -18

OpenStudy (anonymous):

g < 2 g > 2 g < 10 g > 10

OpenStudy (muzzack):

-6g + 3g + 12 > -18 -3g+12>-18 -3g>-30 g<10

OpenStudy (muzzack):

C is ur answer

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

-6g + 3g + 12 = -18 -3g + 12 = -18 -3g = -18 - 12 -3g = -30 g = 10